OpenStudy (anonymous):

i need help solving this problem 2xsquared plus 3xy - 4ysquared when x=2 and y=negative 4

OpenStudy (anonymous):

2x^2 + 3xy - 4y^2 x= 2 y = -4 2*(2)^2 + 3 *2*(-4) -4(-4)^2 = 8 -24 -4*16= -16 -64 = - 80